UT 20131014UT FLAO+PISCES DX

Observer: Arcidiacono,C, Boutsia, K, (INAF), D.Miller (LBTO)
Telescope Operator:Bechetti G
SW Support:
AO Support: D.Miller
Telescope Support: Kevin
Instrument Support:

Summary

We were open from 5:00UT - 7:00UT but we did not succeed in taking any useful data - bad seeing 1.5"-3.0"

Project_RUN Objects Filter ExpTime Notes
Antoniucci PVCep H2,FeII,K 1s each we had a first sucessful preset but after that clouds/wind/humidity came in. We could not apply AO corrections since the ref star was R=16.1 (yesterday R=15.3) - at the end of the preset it was too faint to be measured - probable clouds and bad seeing - star was too faint for the ccd47: the bias on ccd47 had to be fine tuned to actually get the star

Details

05:06 The dome is open - the TO checks pointing and collimation

05:27 Slew to PVCep
  • first preset failed because the ccd47 bias was not correct/optimal

05:33 send preset again
  • we took some test images in the 4 filters - it does not look saturated

05:51 the shell ripped
  • send preset again
  • there are problems collimating

06:01 send preset again
  • seeing: GCS 2.5" - DIMM 2.2"
  • the guide star seems to be too faint - we choose a brighter star

06:10 home the AGW because the info is correct but it cannot find the star
  • fail again - restart GCS
  • something seems to be failing with the guiding - it keeps failing with different stars

06:20 the catalog does not correspond to the image - find a star from catalog and preset again
  • ccd47 error - wrong bias
  • preset again - the bias has to be set manually on the ccd47
  • the AOref star is too faint
  • we give up

06:31 we slew to a bright star in the North to test the AO system
  • there might be vignetting on the AO guider
  • the seeing has blown up - the wind is picking up
  • preset failed - the AOS is in the safe skip mode - error

06:49 AO failed - we don't know why...
  • preset again
  • test AO with bin1 - it works - we also try a dither pattern to check if everything is working and we will probably close down after that since the wind is picking up and humidity also

06:56 The TO is closing the dome!!

-- KonstantinaBoutsia - 14 Oct 2013
Topic revision: r2 - 14 Oct 2013, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback